For a polutropic process
 
PV^n = constant
 
For what value of n, the polytropic process become isochoric? Explain with reason.

Grade:11

3 Answers

Arun
5 years ago
Dear student
 
we can alsowrite it as
 
P1/n V = c1/n
 
now as we know, isochoric process means volume is constant.
 
hence n should be infinity
ABC
5 years ago
Cannot understand it, can u explain in more detail?
...................................
...................................
Arun
5 years ago
Dear student
 
polytropic process is one in which Volume is constant.
or we can say V = constant
 
But we are given with PVn = C  (where C is a constant)
 
 
Now we have to think that how we can get V = constant
 
hence divide by 1/n in powers
 
P1/n V = C1/n
 
we haveqve to eliminate the term of P
 
hence put n = infinity
 
which equals to V = constant
 
hence n = infinity for isochpric process